Pair equations, one representing a supply curve and other demand curve.

a) Determine coordinates of the market equilibrium b) determine the revenue equation c) determine the revenue at market equilibrium.
Please explain how did you get the answer!!! Thanks
2p+x-300=0, 8p-x-200=0
1
Expert's answer
2013-02-27T06:24:09-0500
a) We need solve the system of two equations to find equilibrium price (p) and quantity (x). Let's do it by adding them.
/2p+x-300=0
+
\8p-x-200=0
& 10p - 500 = 0
p = 50
x = 300 - 2*50 = 200
b) The revenue equation is TR = p*x
c) TR = p*x = 50*200 = 1000